# DMS*: Minimizing Makespan for Multi-Agent Combinatorial Path Finding

### v1

@@ Abstract
Multi-Agent Combinatorial Path Finding (MCPF) seeks collision-free paths for multiple agents from their initial to goal locations, while visiting a set of intermediate target locations in the middle of the paths. MCPF is challenging as it involves both planning collision-free paths for multiple agents and target sequencing, i.e., solving traveling salesman problems to assign targets to and find the visiting order for the agents. Recent work develops methods to address MCPF while minimizing the sum of individual arrival times at goals. Such a problem formulation may result in paths with different arrival times and lead to a long makespan, the maximum arrival time, among the agents. This paper proposes a min-max variant of MCPF, denoted as MCPF-max, that minimizes the makespan of the agents. While the existing methods (such as MS*) for MCPF can be adapted to solve MCPF-max, we further develop two new techniques based on MS* to defer the expensive target sequencing during planning to expedite the overall computation. We analyze the properties of the resulting algorithm Deferred MS* (DMS*), and test DMS* with up to 20 agents and 80 targets. We demonstrate the use of DMS* on differential-drive robots.
Index Terms:
Path Planning for Multiple Mobile Robots or Agents, Multi-Agent Path Finding, Traveling Salesman Problem
@@ I Introduction
Multi-Agent Path Finding (MAPF) seeks a set of collision-free paths for multiple agents from their respective start to goal locations. This paper considers a generalization of MAPF called Multi-Agent Combinatorial Path Finding (MCPF), where the agents need to visit a pre-specified set of intermediate target locations before reaching their goals. MAPF and MCPF arise in applications such as surveillance [1] and logistics [2]. For instance, factories [3, 4] use a fleet of mobile robots to visit a set of target locations to load machines for manufacturing. These robots share a cluttered environment and follow collision-free paths. In such settings, MAPF problems and their generalizations naturally arise to optimize operations.
@@
MCPF is challenging as it involves both collision avoidance among the agents as present in MAPF, and target sequencing, i.e., solving Traveling Salesman Problems (TSPs) [5, 6] to specify the allocation and visiting orders of targets for all agents. Both the TSP and the MAPF are NP-hard to solve to optimality [7, 5], and so is MCPF. A few methods have been developed [8, 9] to address the challenges in MCPF, and they often formulate the problem as a min-sum optimization problem, denoted as MCPF-sum, where the objective is to minimize the sum of individual arrival times. Such a formulation may result in an ensemble of paths where some agents arrive early while others arrive late, which leads to long execution times before all agents finish their paths. This paper thus proposes a min-max variant of MCPF (Fig. 1), denoted as MCPF-max, where the objective is to minimize the maximum arrival time, which is also called the makespan, of all agents.
To solve MCPF-max, this paper first adapts our prior MS* algorithm [8], which was first designed for MCPF-sum, to address MCPF-max. Then, we further develop two new techniques to expedite the planning, and we call the resulting new algorithm Deferred MS* (DMS*).
@@
Figure 1: MCPF-max problem, which seeks a set of collision-free paths while minimizing the maximum arrival time of all agents. The color of a target or goal indicates the assignment constraints, i.e., the subset of agents that are eligible to visit that target or goal. The orange target appears on the path of the blue agent, but is not claimed to be visited by the blue agent.
@@
Specifically, the existing MS* follows the general heuristic search framework (such as A*) by iteratively generating, selecting and expanding states to construct partial solution paths from the initial state to the goal state. MS* uses Traveling Salesman Problem (TSP) algorithms to compute target sequences for agents. When solving the TSP, agent-agent collision are ignored, and the cost of resulting target sequences are thus lower bounds of the true costs for the agents to arrive at their goals. Therefore, the cost of target sequences provides an admissible heuristic to guide the state selection and expansion as in A*. Furthermore, MS* leverages the idea in M* [10] to first use the target sequences to build a low-dimensional search space, and then grow this search space by coupling agents together for planning only when collision happens. By doing so, MS* interleaves TSP (target sequencing) and MAPF (collision resolution) techniques under the heuristic search framework and is able to provide completeness and solution quality guarantees.
@@
The second technique developed in this paper is only applicable to MS* when solving MCPF-max, and does not work for MCPF-sum. Since the goal here is to minimize the makespan, during the search, agents with non-maximum arrival time naturally have “margins” in a sense that they can take a longer path without worsening the makespan of all agents. We take advantage of these margins to let agents re-use their previously computed target sequences and defer the expensive calls of TSP solvers until the margin depletes.
To verify the methods, we conduct both simulation in various maps with up to 20 agents and 80 targets, as well as real robot experiments. The simulation shows that the new techniques in DMS* help triple the success rates and reduce the average runtime to solution. The real robot experiments demonstrate that the planned path are executable on real robots, and that DMS* is able to optimize makespan while the existing planners that minimize the sum of arrival times can lead to long makespan.
@@ II Related Work
II-A Multi-Agent Path Finding
To find collision-free paths for multiple agents, a variety of MAPF algorithms were developed, falling on a spectrum from coupled [11] to decoupled [12], trading off completeness and optimality for scalability. In the middle of this spectrum lie the dynamically-coupled methods such as M* [10] and CBS [13], which begin by planning for each agent a shortest path from the start to the goal ignoring any potential collision with the other agents, and then couple agents for planning only when necessary in order to resolve agent-agent collision. These dynamically-coupled methods were improved and extended in many ways [14, 15, 16, 17].
II-B Traveling Salesman Problems
In the presence of multiple intermediate targets, the planner needs to determine both the assignment and visiting order of the targets for the agents. For a single agent, the well-known Traveling Salesman Problem (TSP), which seeks a shortest tour that visits every vertex in a graph, is one of the most well-known NP-hard problems [5]. Closely related to TSP, the Hamiltonian Path Problem (HPP) requires finding a shortest path that visits each vertex in the graph from a start vertex to a goal vertex. The multi-agent version of the TSP and HPP (denoted as mTSP and mHPP, respectively) are more challenging since the vertices in the graph must be allocated to each agent in addition to finding the optimal visiting order of vertices. To simplify the presentation, we refer to all these problems simply as TSPs. To solve TSPs, a variety of methods have been developed ranging from exact methods (branch and bound, branch and price) [5] to heuristics [18] and approximation algorithms [19], trading off solution optimality for runtime efficiency. This work does not develop new TSPs solvers and leverage the existing ones.
II-C Target Assignment, Sequencing and Path Finding
Some recent research combines MAPF with target assignment and ordering [20, 21, 22, 23, 24, 25, 9, 8, 26, 27, 28, 29]. Most of them either consider target assignment only (without the need for computing visiting orders of targets) [20, 21, 22], or consider the visiting order only given that each agent is pre-allocated a set of targets [23, 24, 25]. Our prior work [8, 9] seeks to handle the challenge in target assignment and ordering as well as the collision avoidance in MAPF simultaneously. These work uses the MCPF-sum formulation and the developed planners minimize the sum of individual arrival times, while this paper investigates MCPF-max, which seeks to minimize the maximum arrival time.
@@ III Problem
Let index set $I=\{1,2,\dots,N\}$ denote a set of $N$ agents. All agents share a workspace that is represented as an undirected graph $G^{W}=(V^{W},E^{W},c^{W})$, where $W$ stands for workspace. Each vertex $v\in V^{W}$ represents a possible location of an agent. Each edge $e=(u,v)\in E^{W}\subseteq V^{W}\times V^{W}$ represents an action that moves an agent between $u$ and $v$. $c^{W}:E^{W}\rightarrow(0,\infty)$ maps an edge to its positive cost value. In this paper, the cost of an edge is equal to its traversal time, and each edge has a unit cost. [^1] In [8, 9], the edge cost can be different from its traversal time: the edge cost can be any positive number while the traversal time is always one unit per edge. In this paper, the edge cost are the same as the edge traversal time, which is one for each edge.
Let the superscript $i\in I$ over a variable denote the specific agent to which the variable belongs (e.g. $v^{i}\in V^{W}$ means a vertex corresponding to agent $i$). Let $v_{o}^{i},v_{d}^{i}\in V^{W}$ denote the initial (or original) vertex and the goal (or destination) vertex of agent $i$ respectively. Let $V_{o},V_{d}\subset V^{W}$ denote the set of all initial and goal vertices of the agents respectively, and let $V_{t}\subset V^{W}$ denote the set of target vertices. For each target $v\in V_{t}$, let $f_{A}(v)\subseteq I$ denote the subset of agents that are eligible to visit $v$; these sets are used to formulate the (agent-target) assignment constraints.
@@
Definition 1 (MCPF-max Problem).
The Multi-Agent Combinatorial Path Finding with Min-Max Objective (MCPF-max) seeks to find a set of conflict-free paths for the agents such that (1) each target $v\in V_{t}$ is visited [^2] In MCPF-max, the notion that an agent $i$ “visits” a target $v\in V_{t}$ means (i) there exists a time $t$ such that agent $i$ occupies $v$ along its path, and (ii) the agent $i$ claims that $v$ is visited. In other words, if a target $v$ is in the middle of the path of agent $i$ and agent $i$ does not claim $v$ is visited, then $v$ is not considered as visited. Additionally, a visited target $v$ can appear in the path of another agent. For the rest of the paper, when we say an agent or a path “visits” a target, we always mean the agent “visits and claims” the target. at least once by some agent in $f_{A}(v)$, (2) the path for each agent $i\in I$ starts at its initial vertex and terminates at a unique goal vertex $v^{i}_{d}\in V_{d}$ such that $i\in f_{A}(v^{i}_{d})$, and (3) the maximum of the cost of all agents’ paths (i.e., $\max_{i\in I}cost(\pi^{i}(v^{i}_{o},v^{i}_{d})))$ reaches the minimum.
Remark 1.
The MCPF-max is similar to MCPF-sum [9, 8]. The only difference between these two problems is that: MCPF-sum seeks to minimize the sum of individual path costs while MCPF-max minimizes the maximum of individual path costs.
@@ IV Method
This section begins with a toy example in Fig. 2 to illustrate the planning process of DMS*, and provides the explanation in the caption. This section then introduces the concepts and notations, explains the pseudo-code and elaborates the technical detail.
Figure 2: An illustration of DMS* and related concepts. (a) shows the workspace graph $G^{W}$ and a toy problem. (b) shows the corresponding target graph $G^{T}$ where each edge in $G^{T}$ corresponds to a minimum cost path in $G^{W}$ between the respective vertices. (c) shows the workflow of DMS* in Alg. 1. (d) shows that DMS* first ignores any agent-agent conflict and solves a corresponding mHPP, which provides a joint sequence $\gamma(l_{o})$ (d1). This joint sequence $\gamma(l_{o})$ can be converted to a joint path $\pi(l_{o})$ (d2,d3), whose corresponding makespan $f_{max}$ is 9. (e) The joint path $\pi(l_{o})$ leads to a policy $\phi_{\gamma(l_{o})}$ that maps a joint vertex to the next joint vertex along $\pi(l_{o})$. Since agent-agent conflicts are ignored, when following the policy, agents may run into conflict (e2). When a conflict is detected, the subset of agents that are in conflict $I_{C}(l)$ is back propagated to the nearby ancestor labels. (f) After the back propagation, these ancestor labels and re-opened and re-expanded while considering all possible actions of the agents that are in conflict (f1,f2,f3). During this re-expansion, for each of the successor, DMS* first uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic of the generated labels to avoid solving a mHPP. For both labels (f2) and (f3), this rough heuristic is $(9,5)-(1,1)=(8,4)$, where $(9,5)$ is the heuristic of (f1), the parent of (f2) and (f3), and $(1,1)$ means that each agent can move at most one step towards their goals. When either of (f2) and (f3) is popped from OPEN and before being expanded, DMS* re-computes a new heuristic and see if the popped label should be expanded or re-added to OPEN for future expansion. To obtain this new heuristic, DMS* first attempts to let the agents follow the previously computed target sequences of the parent label and check if this will worsen the overall makespan. For the successor shown in (f2), the resulting makespan is 9 (f4), which is no worse than the previous makespan 9, and DMS* will not call mHPP solver to save computational effort. For the successor shown in (f3), the resulting makespan is 10 (f5), which is worse than the previous makespan 9, and DMS* have to call mHPP solver to find an optimal joint sequence from that successor (f3) to the goals. Finally, with the joint sequence, new policy for the successors can be built and the search continues as in (e).
@@ IV-A1 Joint Graph
Let $\mathcal{G^{W}}=(\mathcal{V^{W}},\mathcal{E^{W}})=\underbrace{G^{W}\times G^{W}\times\dots\times G^{W}}_{\text{$N$ times}}$ denote the joint graph of all agents which is the Cartesian product of $N$ copies of $G^{W}$, where each vertex $v\in\mathcal{V^{W}}$ represents a joint vertex and $e\in\mathcal{E^{W}}$ represents a joint edge that connects a pair of joint vertices. The joint vertex $v_{o}=(v^{1}_{o},v^{2}_{o},\cdots,v^{N}_{o})$ corresponding to the initial vertices of all agents is called the initial joint vertex. Let $\pi(u,v),u,v\in\mathcal{V^{W}}$ denote a joint path, which is a tuple of $N$ (individual) paths of the same length, i.e., $\pi(u,v)=(\pi^{1}(u^{1},v^{1}),\cdots,\pi^{N}(u^{N},v^{N}))$. For the rest of this paper, we use a “path” to denote an “individual path”. DMS* searches the joint graph $\mathcal{G^{W}}$ for a joint path that solves the MCPF-max.

### v2

@@ Abstract
Multi-Agent Combinatorial Path Finding (MCPF) seeks collision-free paths for multiple agents from their start to goal locations, while visiting a set of intermediate target locations in the middle of the paths. MCPF is challenging as it involves both planning collision-free paths for multiple agents and target sequencing, i.e., solving traveling salesman problems to assign targets to and find the visiting order for the agents. Recent work develops methods to address MCPF while minimizing the sum of individual arrival times at goals. Such a problem formulation may result in paths with different arrival times and lead to a long makespan, the maximum arrival time, among the agents. This paper proposes a min-max variant of MCPF, denoted as MCPF-max, that minimizes the makespan of the agents. While the existing methods (such as MS*) for MCPF can be adapted to solve MCPF-max, we further develop two new techniques based on MS* to defer the expensive target sequencing during planning to expedite the overall computation. We analyze the properties of the resulting algorithm Deferred MS* (DMS*), and test DMS* with up to 20 agents and 80 targets. We demonstrate the use of DMS* on differential-drive robots.
@@ I Introduction
Multi-Agent Path Finding (MAPF) seeks a set of collision-free paths for multiple agents from their respective start to goal locations. This paper considers a generalization of MAPF called Multi-Agent Combinatorial Path Finding (MCPF), where the agents need to visit a pre-specified set of intermediate target locations before reaching their goals. MAPF and MCPF arise in applications such as logistics [1]. For instance, factories use a fleet of mobile robots to visit a set of target locations to load machines for manufacturing. These robots share a cluttered environment and follow collision-free paths. In such settings, MAPF problems and their generalizations naturally arise to optimize operations.
@@
MCPF is challenging as it involves both collision avoidance among the agents as in MAPF, and target sequencing, i.e., solving Traveling Salesman Problems (TSPs) [2, 3] to specify the assignment and visiting orders of targets for all agents. Both the TSP and the MAPF are NP-hard to solve to optimality [4, 2], and so is MCPF. A few methods were developed [5, 6] to address MCPF, and they often formulate the problem as a min-sum optimization problem, denoted as MCPF-sum, where the objective is to minimize the sum of individual arrival times. Such a formulation may result in an ensemble of paths where some agents arrive early while others arrive late, which leads to long execution times before all agents finish their paths. This paper thus proposes a min-max variant of MCPF (Fig. 1), denoted as MCPF-max, where the objective is to minimize the maximum arrival time, which is also called the makespan, of all agents.
@@
Figure 1: MCPF-max and MCPF-sum. MCPF-max seeks a set of collision-free paths while minimizing the maximum arrival time of the agents. The color of a target or goal indicates the assignment constraints, i.e., the subset of agents that are eligible to visit that target or goal.
@@
To solve MCPF-max, we first adapt our prior MS* algorithm [5], which was designed for MCPF-sum, to address MCPF-max. Then, we further develop two new techniques to expedite the planning, and we call the resulting new algorithm Deferred MS* (DMS*). Specifically, the existing MS* is a heuristic search approach (such as A*) by iteratively generating, selecting and expanding states to construct partial solution paths from the initial state to the goal state. MS* uses Traveling Salesman Problem (TSP) algorithms to compute target sequences for the agents. When solving the TSP, agent-agent collision are ignored, and the cost of resulting target sequences are thus lower bounds of the true costs to reach the goals. Therefore, the cost of target sequences provides an admissible heuristic to guide the state selection and expansion as in A*. Furthermore, MS* leverages the idea in M* [7] to first use the target sequences to build a low-dimensional search space, and then grow this search space by coupling agents together for planning only when collision happens. By doing so, MS* interleaves TSP (target sequencing) and MAPF (collision resolution) techniques using a heuristic search approach, and provides completeness and solution optimality.
@@
The second technique is only applicable to MS* when solving MCPF-max, and does not work for MCPF-sum. Since the goal here is to minimize the makespan, during the search, agents with non-maximum arrival time naturally have “margins” in a sense that they can take a longer path without worsening the makespan of all agents. We take advantage of these margins to let agents re-use their previously computed target sequences and defer the expensive calls of TSP solvers until the margin depletes.
To verify the methods, we conduct both simulation in various maps with up to 20 agents and 80 targets, as well as a simple real robot experiment. The simulation shows that: (i) DMS* finds paths that are up to 50% cheaper than an iterative greedy baseline method, and (ii) the new techniques in DMS* help triple the success rates and reduce the average runtime to solution comparing to MS*. The robot experiments show that the planned path are executable, and inspire future work.
@@ II Related Work
Multi-Agent Path Finding algorithms fall on a spectrum from coupled [8] to decoupled [9], trading off completeness and optimality for scalability. In the middle of this spectrum lie the dynamically-coupled methods such as M* [7] and CBS [10], which begin by planning for each agent a shortest path from the start to the goal ignoring any potential collision with the other agents, and then couple agents for planning only when necessary to avoid agent-agent collision.
Traveling Salesman Problems determine both the assignment and visiting order of the targets for the agents, where there are multiple intermediate targets to visit. For a single agent, the Traveling Salesman Problem (TSP) seeks a shortest tour that visits every vertex in a graph, and is a well-known NP-hard problem [2]. Closely related to TSP, the Hamiltonian Path Problem (HPP) requires finding a shortest path that visits each vertex in the graph from a start vertex to a goal vertex. The multi-agent version of the TSP and HPP (denoted as mTSP and mHPP, respectively) are more challenging since the vertices in the graph must be allocated to each agent in addition to finding the optimal visiting order of vertices. We refer to all these problems simply as TSPs. Different methods have been developed [2, 11, 12] to solve TSPs, trading off solution optimality for runtime efficiency. This paper does not develop new TSPs solvers and leverage the existing ones.
Target Assignment, Sequencing and Path Finding were recently combined in different ways [13, 14, 15, 16, 6, 5, 17, 18]. Most of them either consider target assignment only (without the need for computing visiting orders of targets) [13, 14, 15], or consider the visiting order only given that each agent is pre-allocated a set of targets [16, 17, 18]. Our prior work [5, 6] seeks to handle the challenge in target assignment and ordering and the collision avoidance in MAPF simultaneously. These work uses the MCPF-sum formulation and the developed planners minimize the sum of individual arrival times, while this paper investigates MCPF-max. We do not extend our prior method CBSS [6] for MCPF-sum to MCPF-max, because CBSS requires solving a min-sum K-best sequencing problem, and it is not obvious how to handle the min-max variant of this K-best sequencing problem.
@@ III Problem
Let the index set $I=\{1,2,\dots,N\}$ denote a set of $N$ agents. All agents share a workspace that is represented as an undirected graph $G^{W}=(V^{W},E^{W},c^{W})$, where $W$ stands for workspace. Each vertex $v\in V^{W}$ represents a possible location of an agent. Each edge $e=(u,v)\in E^{W}\subseteq V^{W}\times V^{W}$ represents an action that moves an agent between $u$ and $v$. $c^{W}:E^{W}\rightarrow(0,\infty)$ maps an edge to its positive cost value. In this paper, the cost of an edge is equal to its traversal time, and each edge has a unit cost. [^1] Here, the edge cost is the same as the edge traversal time, which is one unit for each edge. When the traversal time of edges are not unitary, it leads to continuous-time MAPF, and we refer the reader to [19, 20].
Let the superscript $i\in I$ over a variable denote the specific agent to which the variable belongs (e.g. $v^{i}\in V^{W}$ means a vertex corresponding to agent $i$). Let $v_{o}^{i},v_{d}^{i}\in V^{W}$ denote the initial (or original) vertex and the goal (or destination) vertex of agent $i$ respectively. Let $V_{o},V_{d}\subset V^{W}$ denote the set of all initial and goal vertices of the agents respectively, and let $V_{t}\subset V^{W}\backslash\{V_{o}\cup V_{d}\}$ denote the set of target vertices. For each target $v\in V_{t}$, let $f_{A}(v)\subseteq I$ denote the subset of agents that are eligible to visit $v$; these sets are used to formulate the (agent-target) assignment constraints.222 An agent $i$ “visits” a target $v\in V_{t}$ means (i) there exists a time $t$ such that agent $i$ occupies $v$ along its path, and (ii) the agent $i$ claims that $v$ is visited. If a target $v$ is in the middle of the path of $i$ and $i$ does not claim $v$ is visited, then $v$ is not considered as visited. A visited target $v$ can appear in the path of another agent. When we say an agent or a path “visits” a target, we always mean the agent “visits and claims” the target. The assignment constraints do not forbid any agent $j\notin f_{A}(v)$ to use $v$ in its path, and only forbid agent $j$ to claim visiting $v$.
@@
Definition 1 (MCPF-max Problem)
The MCPF with Min-Max Objective (MCPF-max) seeks to find a set of conflict-free paths for the agents such that (1) each target $v\in V_{t}$ is visited at least once by some agent in $f_{A}(v)$, (2) the path for each agent $i\in I$ starts at its initial vertex and terminates at a unique goal vertex $v^{i}_{d}\in V_{d}$ such that $i\in f_{A}(v^{i}_{d})$, and (3) the maximum of the cost of all agents’ paths (i.e., $\max_{i\in I}g(\pi^{i}(v^{i}_{o},v^{i}_{d})))$ reaches the minimum.
@@ IV Method
This section begins with an example in Fig. 2 to illustrate the planning process of DMS*. Then, Sec. IV-A introduces some concepts that will be used during the search process, before the presentation of the search algorithm in Sec. IV-B.
Figure 2: An illustration of DMS* and related concepts. (a) shows the workspace graph $G^{W}$. (b) shows the target graph $G^{T}$ where each edge in $G^{T}$ corresponds to a minimum cost path in $G^{W}$ between the respective vertices. (c) shows the workflow of DMS* in Alg. 1. (d) shows DMS* first ignores any agent-agent conflict and solves a corresponding mHPP, which provides a joint sequence $\gamma(l_{o})$ (d1). This joint sequence $\gamma(l_{o})$ can be converted to a joint path $\pi(l_{o})$ (d2,d3), whose corresponding makespan $f_{max}$ is 9. (e) The joint path $\pi(l_{o})$ leads to a policy $\phi_{\gamma(l_{o})}$ that maps a joint vertex to the next joint vertex along $\pi(l_{o})$. Since conflicts are ignored in this policy, agents may run into conflict (e2). When a conflict is detected, the subset of agents that are in conflict $I_{C}(l)$ is back propagated to the ancestor labels. (f) After the back propagation, these ancestor labels and re-opened and re-expanded while considering all possible actions of all agents in conflict (f1,f2,f3). During the re-expansion, for each of the successors, DMS* first uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic of the generated labels to avoid solving a mHPP. For both labels (f2) and (f3), this rough heuristic is $(9,5)-(1,1)=(8,4)$, where $(9,5)$ is the heuristic of (f1), the parent of (f2) and (f3), and $(1,1)$ means each agent can move at most one step towards their goals. When either (f2) or (f3) is popped from OPEN and before being expanded, DMS* re-computes a new heuristic and checks if the popped label should be expanded or re-added to OPEN for future expansion. To obtain this new heuristic, DMS* first attempts to let the agents follow the previously computed target sequences of the parent label and check if this worsens the makespan. For the successor shown in (f2), the resulting makespan is 9 (f4), which is no worse than the previous makespan 9, and DMS* will not call mHPP solver to save computational effort. For the successor shown in (f3), the resulting makespan is 10 (f5), which is worse than the previous makespan 9, and DMS* have to call mHPP solver to find a new joint sequence from that successor (f3) to the goals. Finally, with the joint sequence, new policy for the successors can be built and the search continues as in (e).
@@ IV-A1 Joint Graph
Let $\mathcal{G^{W}}=(\mathcal{V^{W}},\mathcal{E^{W}})=\underbrace{G^{W}\times G^{W}\times\dots\times G^{W}}_{\text{$N$ times}}$ denote the joint graph of the agents which is the Cartesian product of $N$ copies of $G^{W}$, where each $v\in\mathcal{V^{W}}$ represents a joint vertex and $e\in\mathcal{E^{W}}$ represents a joint edge that connects a pair of joint vertices. Let $v_{o}=(v^{1}_{o},v^{2}_{o},\cdots,v^{N}_{o})$ denote the initial joint vertex, which contains the initial vertices of all the agents. Let $\pi(u,v),u,v\in\mathcal{V^{W}}$ denote a joint path, which is a tuple of $N$ (individual) paths of the same length, i.e., $\pi(u,v)=(\pi^{1}(u^{1},v^{1}),\cdots,\pi^{N}(u^{N},v^{N}))$. A joint path $\pi$ can be viewed as a list of joint vertices $(v_{0},v_{1},\cdots,v_{\ell})$, and the time for the agents to arrive at each joint vertex $v\in\pi$ is specified by the index of $v$ in $\pi$. Given two subsequent joint vertices $u,v$ along a joint path $\pi$, let $\textsl{CheckConflict}(u,v)$ denote a procedure that checks for vertex conflicts at $u,v$ and edge conflicts during the transition from $u$ to $v$ among all pairs of agents, and CheckConflict returns a set $I_{C}\subseteq I$ of agents that are in conflict. We use a “path” to denote an “individual path”. DMS* searches the joint graph $\mathcal{G^{W}}$ for a joint path that solves the MCPF-max.